OpenStudy (anonymous):

A rock falls 30 m. Its KE before it strikes the ground is 4,392 J. What is the mass of the rock?

OpenStudy (shamim):

Given h=30m K.E=4392J m=?

OpenStudy (shamim):

At 30m height the potential energy is P.E=mgh

OpenStudy (shamim):

Before touching the ground kinetic energy is K.E=mgh 4392=m*9.8*30

OpenStudy (shamim):

m=?

OpenStudy (shamim):

Response plz!!"

OpenStudy (anonymous):

146kg

OpenStudy (shamim):

No

OpenStudy (shamim):

m=4392/(9.8*30)=?